S6. Remember Horatio? Once you convinced him that he should not apply for the Nobel Prize, he turned his attention to his hobby of stretching rubber bands. One day he accidentally touched a freshly stretched rubber band to his cleanly shaven face (Horatio, believed in obeying the dress and grooming code) and observed that heat was being evolved when the band was stretched. Excited by this discovery he raced to the lab and performed the following experiments. (1) A rubber band was stretched at room temperature and the length recorded as 11. The rubber band was then heated and the new length, 12, recorded (2) Horatio then calorimetrically measured Q by reversibly stretching the rubber band (impossible) from 10 to 210. During this reversible constant-temperature (27°C) process, 6.28 J of heat was evolved. The rubber band was then allowed to contract irreversible (the way nature intended a rubber band to contract) back to 10 without any restraining tension. The heat absorbed was 2.09 J a. Is 12 longer or shorter than 11 Explain. b. During step 2, how much work was required to stretch the rubber band? c. Calculate the entropy change of the rubber band for stretching from 10 to 210o d. Calculate the change in entropy for the contraction of the rubber band e. What is the change in entropy of the universe for the two-step process of stretching and contracting the rubber band?
Answer #1

a)

L2 is longer. Because it has absorbed some heat already and possesses more energy

b)

6.28 J. Same as that of heat released

c)

s = Q/T = 6.28 / (27+273) = 6.28/300 = 0.021 J/K

d)

s = Q/T = 2.09/(300 = 0.007 J/K

e)

Zero. As reversible processes show no change in entropy